Dillon Labs has asked its financial manager to measure the cost of each specific type of capital as well as the weighted average cost of capital. The weighted average cost is to be measured by using the following​ weights:

40​%

​long-term debt,

25​%

preferred​ stock, and

35​%

common stock equity​ (retained earnings, new common​ stock, or​ both). The​ firm's tax rate is

22​%.

Debt The firm can sell for

​$1020

a

20​-year,

​$1,000​-par-value

bond paying annual interest at a

8.00​%

coupon rate. A flotation cost of

3​%

of the par value is required.

Preferred stock  

9.00​%

​(annual dividend) preferred stock having a par value of

​$100

can be sold for

​$88

An additional fee of

​$5

per share must be paid to the underwriters.

Common stock  The​ firm's common stock is currently selling for

​$80

per share. The stock has paid a dividend that has gradually increased for many​ years, rising from

​$2.00

ten years ago to the

​$3.42

dividend​ payment,

Upper D 0D0​,

that the company just recently made. If the company wants to issue new new common​ stock, it will sell them

​$3.00

below the current market price to attract​ investors, and the company will pay

​$3.00

per share in flotation costs.

 Calculate the​ after-tax cost of debt.6.32

The​ after-tax cost of debt using the approximation formula is

nothing​ 6.31

Calculate the cost of preferred stock 10.84

 Calculate the cost of common stock​ (both retained earnings and new common​ stock) 10.02 and 10.39

 Calculate the WACC for Dillon Labs.
Answer #1
a) Price of the bond = 1020-1000*3% = $             990
Before tax cost of debt = YTM.
YTM using an online calculator = 8.10%
After tax cost of debt = 8.10%*(1-22%) = 6.32%
b) Cost of preferred stock = 9/(88-5) = 10.84%
c) Growth rate in dividends = (3.42/2.00)^(1/10)-1 = 5.51%
Cost of retained earnings = 3.42*1.0551/80 +0.0551= 10.02%
Cost of new equity = 3.42*1.0551/(80-3-3)+0.0551 = 10.39%
d) WACC:
Using retained earnings = 6.32%*40%+10.84%*25%+10.02%*35% = 8.75%
Using new common stock = 6.32%*40%+10.84%*25%+10.39%*35% = 8.87%